For 50,000, Kelly purchases an annuity-immediate that pays 400 monthly for the next 20 years. Calculate the annual nominal interest rate convertible monthly earned by Kelly's investment.

(Don't use Excel)
Answer #1
Amount of Equal payment annuity= 400
Amount paid for annuity (P)= 50000
Total months in 20 years (n)=20*12= 240

Annuity or Equal payments formula = P* i *((1+i)^n)/(((1+i)^n)-1)

400 = 50000*i*((1+i)^240)/(((1+i)^240)-1)

We will Assume ínterest rate per month is 0.6% or 0.006

50000*0.006*((1+0.006)^240)/(((1+0.006)^240)-1)

393.6746495
We will Assume ínterest rate per month is 0.7% or 0.007

50000*0.007*((1+0.007)^240)/(((1+0.007)^240)-1)

430.7522477

Actual Payment is in between two payments calculated by i. so We will calculate (i) by interpolation formula

interpolation formula = uper rate - (uper rate - lower rate)*(Uper value - actual value)/(uper value - lower value)

0.007 - ((0.007-0.006)*(430.7522477-400)/(430.752247-393.6746495)

0.007- 0.00082940
0.00617060

Per month rate is 0.00617060

So annual Rate is 0.00617060*12= 0.07405

So nominal annual rate is 7.405 or 7.41% convertible semiannually